The following points fall on a parabola. Write a system of equations that you could use to determine the equation for the parabola(you do not have to solve the system)
This is what is in the box
x -2 0 1
y 39 5 9

Answer by scott8148(6628)

the general equation is ___ y = ax^2 + bx + c

three unknowns and three equations

Eq 1 ___ 39 = a(-2)^2 + b(-2) + c

Eq 2 ___ 5 = a(0)^2 + b(0) + c

Eq 3 ___ 9 = a(9)^2 + b(9) + c